OpenStudy (anonymous):

x ^{sqrt{logx}}=10^{8}

OpenStudy (anonymous):

you need to take the log of both sides, do you know how to do that?

OpenStudy (anonymous):

yea so far i got square root of log of x log x = 8

OpenStudy (anonymous):

then square both sides

OpenStudy (anonymous):

to get rid of the root

OpenStudy (anonymous):

\[\sqrt{\log x} \log x = \log 10^8 \]

OpenStudy (anonymous):

well you made it harder on yourself let me explore this in details

OpenStudy (anonymous):

\[\sqrt{\log_{10}(x) }=1/2\log_{10}(x) \]

OpenStudy (anonymous):

do you agree

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now we have

OpenStudy (anonymous):

\[x ^{\frac{ 1 }{ 2 }\log_{10}(x) }=10^{8}\]

OpenStudy (anonymous):

\[\log[x ^{\frac{ 1 }{ 2 }\log_{10}(x) }=10^{8}]\]

OpenStudy (anonymous):

right and then is \[\frac{ 1}{ 2} (\log x )^2 \] ?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

i think now you know how to do the rest

OpenStudy (anonymous):

all you're left with is multiplying both side by 2 then taking the root of both sides

OpenStudy (anonymous):

i get a different answer though , the book says the answer should be 10,000 :/

OpenStudy (anonymous):

then the anti log

OpenStudy (anonymous):

or log inverse

OpenStudy (anonymous):

its not giving me the 10,000 :(

OpenStudy (anonymous):

i got 10 000

OpenStudy (anonymous):

ok i will show you the whole solution

OpenStudy (anonymous):

but be patient because i am new to this editor

OpenStudy (anonymous):

i am new too no worries thank you so much!

OpenStudy (anonymous):

do you agree with this so far

OpenStudy (anonymous):

\frac{ 1}{ 4} (\log x )^2

OpenStudy (anonymous):

why do you square the 8?

OpenStudy (anonymous):

\[\frac{ 1}{ 4} (\log x )^2=8^{2}\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i meant 16

OpenStudy (anonymous):

because i wrote this on a paper

OpenStudy (anonymous):

\[\ (\log x )^2=16\]

OpenStudy (anonymous):

sorry again

OpenStudy (anonymous):

\[\frac{ 1}{ 2} (\log x )^2 = 8\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[\ (\log x )^2=16\]

OpenStudy (anonymous):

because we multiplied both side by 2

OpenStudy (anonymous):

then we take the square root of both sides

OpenStudy (anonymous):

that yields to

OpenStudy (anonymous):

\[ (\log x )=4\]

OpenStudy (anonymous):

got it !!! omg you are awesome thankss !!!!!!

OpenStudy (anonymous):

you are welcome sorry for the mistakes its about 6 am my time

OpenStudy (anonymous):

omg ! go to sleep!

OpenStudy (anonymous):

where are you from?

OpenStudy (anonymous):

UK

OpenStudy (anonymous):

wow! well thanks again for all your help!

OpenStudy (anonymous):

you are more than welcome at any time

OpenStudy (anonymous):

well good luck :)

OpenStudy (anonymous):

you too